OpenStudy (anonymous):

what is the latus rectum of the parabola x^2-4x-2y-8=0 options are:>> a)8 b)4 c)2 d)1

OpenStudy (anonymous):

For this question we will need to use imaginary numbers, we will need to take the first number and multiply it by eleventeen, then divide the second number by seventytwelve. Then we cross multiply the entire thing and then add fortyfifteen. This will bring you to your answer.

OpenStudy (anonymous):

Separate x values from non- x values. \[x ^{2} −4x=2y+8 \] complete the square for the x side \[(x−2)^{2} =2y+8+4 \] combine and factor (GCF) the right side \[(x−2)^{2} =2(y+6)\] The coefficient of the non-squared side is the length of the latus rectum.